OpenStudy (anonymous):

Find the leg of each isosceles right triangle when the hypotenuse is of the given measure. Given = 5sqrt6

OpenStudy (jdoe0001):

hmmm is it a right triangle?

OpenStudy (anonymous):

\[2x ^{2}=(5\sqrt{6})2\]

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

* \[(5\sqrt{6})^{2}\]

OpenStudy (anonymous):

I got that from Pythagorean theorem. a^2 + b^2 = c^2

OpenStudy (anonymous):

since a and b have to be equal, i just made it 2x^2=c^2

OpenStudy (anonymous):

where c is 5sqrt6

OpenStudy (anonymous):

Ok, so solving for x?

OpenStudy (anonymous):

When i solved for x i got 125?

OpenStudy (anonymous):

did you take the square root?

OpenStudy (anonymous):

I got 5sqrt3 or 8.66

OpenStudy (jdoe0001):

so... hmmm is it a right triangle?

OpenStudy (jdoe0001):

http://www.mathsisfun.com/definitions/hypotenuse.html <--- notice that so... what do you think, is it a right triangle or maybe not?

OpenStudy (jdoe0001):

ohhh wait a sec... seems to have been added now... ohhh alrite

OpenStudy (jdoe0001):

and "isosceles" triangle means two sides are equal a hypotenuse is the "longest" side, so it cannot by definition have another equal so the other two sides should be equal then

OpenStudy (anonymous):

Ok wow i feel dumb i was using the wrong formula the entire time

OpenStudy (anonymous):

that's why my answer was far off

OpenStudy (anonymous):

Did you get it though?

OpenStudy (anonymous):

Yea i got it finally.

OpenStudy (anonymous):

Another question though, Find the diagonal of a square whose sides are of the given measure. Given = 3^2?

OpenStudy (anonymous):

i cant seem to plug this one in right

OpenStudy (anonymous):

given = 3sqrt2*

OpenStudy (anonymous):



OpenStudy (jdoe0001):

 so you could use the pythagorean theorem for that since \(\bf c^2=a^2+b^2\implies \qquad \begin{cases} c=hypotenuse\\ a=adjacent\\ b=opposite \end{cases} \\ \quad \\ c^2=b^2+b^2\qquad \begin{cases} \textit{recall, sides are equal, thus}\\ a=b \end{cases} \\ \quad \\ c^2=2b^2\implies \cfrac{c^2}{2}=b^2\implies \sqrt{\cfrac{c^2}{2}}=b=a \\ \quad \\ \sqrt{\cfrac{{\color{brown}{ (5\sqrt{6}}})^2}{2}}=b=a\)